OpenStudy (anonymous):

WILL FAN AND MEDAL!!! Given the following system of equations, identify the type of system. y = -x + 4 2x + 2y = 8 independent inconsistent equivalent

OpenStudy (tennis5518):

What do you think?

OpenStudy (anonymous):

Inconsistent?

OpenStudy (welshfella):

multiply the first equation by 2 then add 2x to both sides what do you get?

OpenStudy (anonymous):

Equivalent?

OpenStudy (tennis5518):

If one (or more) of the equations can be expressed as a linear combination of the others. This is equivalent to the statements the matrix of coefficients does not have an inverse (or is singular), or the determinant of the matrix of coefficients is zero.

OpenStudy (anonymous):

So it's equivalent???

OpenStudy (anonymous):

@Tennis5518 ^

OpenStudy (tennis5518):

No

OpenStudy (tennis5518):

Wait

OpenStudy (welshfella):

y = -x + 4 * 2 gives 2y = -2x + 8

OpenStudy (welshfella):

now add 2x to both sides and compare it with the second equation

OpenStudy (anonymous):

x=−y+4 @welshfella

OpenStudy (anonymous):

So it's...independent??? I'm still lost

OpenStudy (anonymous):

I finally got a hold of my teacher and we figured out the answer is equivalent :)